# Flowr.root – A flow matching based foundation model for joint multi-purpose structure-aware 3D ligand generation and affinity prediction

### v3

@@ 6 Conclusion
Appendix
Appendix A Datasets
@@
In Sec. 3 in Fig. 1 we visualize the data generation framework that we established for this work. As outlined in Sec. 3 we used as protein-ligand complex data Plinder, BindingMOAD, SPINDR, HiQBind, SAIR, BindingNet, CrossDocked2020, KIBA-3D, Davis-3D, and Kinodata-3D. After downloading and aggregating the different data sources, we split proteins and ligands, if not already done, into separate files. We save ligands as SDF and proteins as PDB or CIF files. Afterwards, we filter the data and remove systems that classify the ligands either as ions, artifacts or not definable. Then, we use Schrodinger’s LigPrep to prepare the ligands and Schrodinger’s PrepWizard to prepare the protein-ligand complex (see Sec. 3 for more details). Successfully prepared systems are converted into our internal pocket-ligand representation encapsulated as PocketComplex instances for easy handling and data management. We retrieve pockets by cutting around the reference ligands with a cutoff of 7A and, if necessary, trim the pocket to the max size of 800 atoms. Every PocketComplex instance is annotated with comprehensive metadata spanning affinity values (if available), GNINA scores, interaction fingerprints, and selected RDKit features, like molecular weight, logP, TPSA, number of hydrogen acceptors and donors, number of rotatable bonds, number of aliphatic rings and many more. Finally, for all datasets we calculate distribution statistics comprising atom, charge, and bond type distributions as well as bond angles, bond lengths and dihedral angles distributions. Importantly, all datasets are splitted following the Plinder train, validation and test splits.
@@
In Fig. 12, we compared the chemical composition of all protein–ligand datasets used in this work by analyzing the distributions of ligand atom types, protein atom types, and protein residue types. The aggregated distributions (top row) reveal the overall prevalence of each type, while the per-dataset heatmaps (bottom row) illustrate how these distributions vary between datasets. For atom types, a logarithmic scale was used to emphasize less frequent elements. Only the 20 standard amino acids were considered for residue type comparisons, and hydrogens were excluded from atom type analyses to improve interpretability.
@@
To provide a comprehensive overview of the affinity data, Fig. 13 summarizes both the distribution and coverage of affinity types across all datasets. The ECDFs display the range and distribution of affinity values for each affinity type, highlighting differences in spread and modality. The overall affinity type counts, presented on a logarithmic scale, reveal a pronounced imbalance, with IC50 being the most abundant. Dataset-specific distributions are visualized using horizontal violin plots, while horizontal bar plots summarize the number of measurements for each affinity type within each dataset. The majority of affinity values originate from the SAIR and BindingNet datasets. Notably, BindingNet also provides Ki measurements, while HiQBind and BindingMOAD offer a balanced representation of IC50, Ki, and Kd values. HiQBind is the only dataset containing EC50 measurements. In contrast, Davis-3D is exclusively annotated with Kd values. This overview facilitates the identification of data biases and gaps that may influence downstream analyses.
@@
To systematically assess the chemical diversity of ligands across datasets, we analyzed the distributions of several continuous molecular descriptors in Fig. 14, including molecular weight, logP, molar refractivity, and topological polar surface area (TPSA). These features capture key aspects of ligand size, hydrophobicity, polarizability, and polarity, which are relevant for molecular recognition and drug-likeness. The results reveal pronounced differences between datasets: for example, BindingNet and SAIR contain ligands with a broader range of molecular weights and higher TPSA values, while datasets such as HiQBind and BindingMOAD exhibit more constrained distributions. These trends reflect the varying selection criteria and source domains of the datasets, and highlight the importance of considering chemical diversity in benchmarking and model development.
@@
We further characterized the datasets by comparing the distributions of discrete ligand features in Fig. 15, including the number of hydrogen bond acceptors and donors, rotatable bonds, ring systems, and chiral centers. These properties provide insight into molecular complexity, flexibility, and the presence of functional groups relevant for binding interactions. The analysis demonstrates that datasets such as BindingNet and SAIR encompass ligands with higher numbers of rotatable bonds and greater ring system diversity, whereas datasets like Davis-3D and KIBA-3D are more restricted in these aspects. Notably, the number of chiral centers and structural alerts also varies substantially, underscoring differences in stereochemical complexity and potential reactivity. Together, these comparisons elucidate the distinct chemical spaces sampled by each dataset and inform the interpretation of downstream modeling results.
A.1 KIBA-3D
@@
The KIBA dataset was originally introduced by Tang et al. [tang2014] as a large-scale benchmark for kinase–ligand binding affinity prediction, integrating heterogeneous bioactivity readouts into a standardized KIBA score. The original dataset contained chemical compounds, kinase targets and over interaction records. While comprehensive, the original KIBA dataset was highly scattered, due to an imbalance between the number of ligand-kinase pairs and associated measurements. To improve the density and ensure more reliable evaluation, subsequent studies discarded the ligands and targets with less than 10 interaction records. This procedure reduced the dataset to ligands and kinase targets. For this work, we prepared the 3D extension of KIBA (KIBA-3D) based on kinase targets with experimentally solved structures deposited in the PDB. Figure shows examples of our docking results, illustrating that Glide could reproduce near-native binding poses for selected kinase inhibitors with conserved bidentate hydrogen bond pattern in the hinge region (e.g., PDBID:2XB7-CHEMBL461876, glide gscore , PDBID:6LVM-CHEMBL1968590, glide gscore ; see Fig. 16. The KIBA-3D dataset serves as a large-scale, structure-consistent benchmark for evaluating generative and predictive models.
A.1.1 Target protein structure selection for KIBA-3D

### v5

@@ Methods
Acknowledgements
F.M. acknowledges support from Helmholtz Munich. E.S. acknowledges support from Jagiellonian University Medical College.
Author Contributions
J.C. conceived the study; developed the methodology; performed experiments, validation, formal analysis; implemented the software; curated the data; created the visualizations; and wrote the original draft of the manuscript. T.L. assisted with experiments. M.M.G. contributed to study design. E.S. assisted with data preparation. F.M. performed validation, formal analysis; created visualizations and wrote the original draft of the manuscript. D.-A.C. provided supervision. All authors reviewed and edited the manuscript.
Competing Interests
J.C., T.L., M.M.G. and D.-A.C. are employees of Pfizer Inc. The remaining authors declare no competing interests.
Supplementary Information
Methods
Flowr.ui: A graphical user interface for interactive ligand design
Flowr.ui is a browser-based graphical user interface for interactive, structure-aware ligand design using the Flowr.root generative model (Supplementary Fig. 9). The application implements a two-tier client-server architecture: a CPU-bound FastAPI [117] frontend server handles molecular file management, input validation, property computation, interaction detection, and chemical space analysis, while a separate GPU-bound FastAPI worker service loads the Flowr.root model checkpoint and executes ligand generation. This separation enables deployment on HPC clusters where the frontend operates on login or CPU nodes and GPU resources are allocated on demand via SLURM, with automatic release after a configurable idle timeout.
Flowr.ui supports both structure-based (SBDD, conditioned on a protein pocket) and ligand-based (LBDD, molecule-only) design workflows. Seven generation modes are available: (1) de novo generation, optionally conditioned on a reference ligand; (2) substructure inpainting, replacing user-selected atoms while retaining fixed substructures; (3) scaffold hopping, replacing the Murcko scaffold while preserving R-groups; (4) scaffold elaboration, decorating a retained ring scaffold with new functional groups; (5) linker inpainting, regenerating atoms connecting ring systems; (6) core growing, retaining a ring system core and regenerating the remainder; and (7) fragment growing, extending from a placed fragment with a configurable anisotropic Gaussian prior center. Users select atoms for retention or replacement directly in the 3D viewer; inpainting masks are computed in real time using pure RDKit-based [119] algorithms without GPU access.
The frontend renders protein structures as interactive 3D cartoon representations with optional molecular surfaces and binding-site residue highlighting (3Dmol.js [118]), 2D molecular depictions (RDKit.js [119]), and analytical plots (Plotly.js [120]). Gaussian prior clouds representing the spatial sampling distribution are visualized as 3D point clouds with interactive repositioning. Post-generation analysis includes chemical space projections via PCA, -SNE [121], or UMAP [122] of Morgan fingerprints [123] (radius 2, 2048 bits), property distribution plots for 20 molecular descriptors (including molecular weight, LogP, TPSA, hydrogen bond donors and acceptors, rotatable bonds, ring counts, fraction of carbons, and structural alerts), affinity distribution tracking, and protein-ligand interaction visualization (hydrogen bonds and salt bridges) rendered as 3D distance-labeled overlays and 2D interaction diagrams.
Generated ligands undergo configurable post-processing: validity and uniqueness filtering, Tanimoto-based diversity filtering across generation rounds, RDKit [119] or GFN2-xTB [124] geometry optimization, property-based filtering, and optional ADMET model filtering. Results are accumulated over multiple rounds with cross-round diversity enforcement. Binding affinities (IC50, Ki, Kd, EC50) are automatically parsed from uploaded SDF property tags with unit detection and p-value conversion, enabling direct comparison between reference and generated compounds. Flowr.ui further implements an active learning loop: users select promising ligands, trigger LoRA [125] finetuning of the loaded Flowr.root checkpoint on the GPU worker, and regenerate with the adapted model, enabling iterative structure-activity relationship refinement within a single session. All generated ligands can be ranked, selected, and exported as SDF files.
For HPC deployment, the server manages the full SLURM lifecycle—job submission with configurable resource parameters, compute node discovery via squeue, worker health probing, and automatic GPU deallocation. Users access the interface via SSH port forwarding. File transfers between frontend and worker are authenticated using HMAC-based tokens, and the application enforces content security policies, rate limiting, and input sanitization.
Datasets
For our small molecule database, we utilized Zinc3D, PubChem3D, Enamine REAL, and OMol25. For our protein-ligand database, we assembled a comprehensive collection of protein–ligand complex datasets by aggregating and standardizing data from multiple sources, including Plinder, BindingMOAD, SPINDR, HiQBind, SAIR, BindingNet, CrossDocked2020, KIBA-3D, Davis-3D, and Kinodata-3D. As visualized in Supplementary Fig. 10, after rigorous preprocessing, filtering, and preparation of both ligands and proteins, each complex was converted into a unified internal representation and annotated with extensive metadata, such as affinity values, if available, and molecular descriptors. This harmonized dataset enables systematic analysis of chemical composition, structural diversity, and affinity distributions across all included sources, providing a robust foundation for downstream modeling and benchmarking.
Importantly, throughout all protein-ligand datasets we kept a consistent dataset split following the provided Plinder [1] train, validation and test set splits to avoid data leakage as best as possible enabling a stringent downstream evaluation. For more details about the datasets we refer to Supplementary Section Datasets, where we provide an overview of the aforementioned dataset generation, curation, and preprocessing pipeline, and a set of different dataset statistics.
Zinc3D
Zinc3D [2] is a subset of the ZINC20 database containing pre-computed low-energy conformations for commercially available compounds. By providing ready-to-use 3D structures, ZINC3D eliminates the computational overhead of on-the-fly conformer generation, thereby accelerating virtual screening campaigns and structure-based drug discovery workflows. We utilized 646,663,126 molecules with their provided conformations.
PubChem3D
PubChem3D [3] extends the widely-used PubChem database by providing up to 500 computed 3D conformations for millions of bioactive compounds. This resource offers extensive conformational diversity, crucial for training robust generative models. We employed 10 conformations per molecule from approximately 93 million compounds, resulting in 928,649,525 total conformations.
Enamine REAL
The Enamine REAL database comprises commercially available, synthetically accessible compounds widely employed in virtual screening and drug design. Its well-curated chemical space supports generative design strategies focused on drug-likeness and synthetic feasibility. Following Cremer et al. [4], we used a diversity subset of the Enamine REAL database. We employed OpenEye’s Omega software with default parameters to generate up to five conformers per molecule, yielding 111,389,149 conformations.
OMol25
OMol25 [5] is a comprehensive dataset containing over 100 million density functional theory (DFT) calculations at the B97M-V/def2-TZVPD level of theory. The dataset covers systems up to 350 atoms with exceptional chemical and structural diversity across 83 elements. We utilized two subsets: the small molecules collection (21,352,259 structures) comprising recomputed versions of widely-used datasets (ANI-2X, Orbnet Denali, SPICE2, Solvated Protein Fragments, and 30% of GEOM) upgraded to consistent high-level DFT theory, and the biomolecules subset (5,180,233 structures), encompassing protein-ligand interactions derived from BioLiP2 and additional complexes generated through docking drug-like molecules from GEOM, ChEMBL, and ZINC20. The biomolecule dataset features protein pocket environments processed through molecular dynamics simulation with appropriate capping and protonation state sampling.
KIBA-3D
We introduce KIBA-3D, a kinase-focused dataset derived from the KIBA bioactivity dataset [6]. Using Schrödinger’s Glide, we performed exhaustive cross-docking of the KIBA ligand space across 172 kinase targets, generating 333,670 protein–ligand complexes spanning 2,038 unique ligands. This dataset provides a dense bioactivity landscape ideally suited for affinity guided, pocket-conditioned generative modeling and kinase-specific benchmarking applications.
Davis-3D
DAVIS-3D is a structurally augmented version of the original DAVIS kinase bioactivity dataset [7], created as part of the Folding-Docking-Affinity (FDA) framework [8]. The dataset was generated by computationally folding protein structures using ColabFold and determining protein-ligand binding conformations through DiffDock, a deep learning-based docking model. This process transforms the original sequence-based DAVIS dataset into a collection of three-dimensional protein-ligand binding structures, of which we utilized 12,982 complexes with binding affinities.
Kinodata-3D
Kinodata-3D [9] is a curated collection of kinase complexes processed using cross-docking methodologies. We utilized 94,211 complexes from the combined mid- to higher-confidence subsets, which include binding affinity annotations and are preprocessed with a 5Å protein-ligand pocket cutoff.
BindingNet
BindingNet [10, 11] comprises 689,796 modeled protein-ligand binding complexes across 1,794 protein targets. The dataset was constructed using an enhanced template-based modeling workflow that incorporates pharmacophore and molecular shape similarities alongside chemical similarity. Structures are categorized by template matching quality into high-confidence (232,030), moderate-confidence (164,912), and low-confidence (292,813) subsets, all of which we employed with their respective binding affinity annotations.
Plinder
Plinder [1] represents the largest and most comprehensively annotated protein-ligand interaction (PLI) dataset, containing 449,383 PLI systems with over 500 annotations per complex. The dataset encompasses diverse interaction types including multi-ligand systems, oligonucleotides, peptides, and saccharides. Plinder introduced an approach for generating training and evaluation splits that minimizes task-specific leakage while maximizing test set quality. After removing complexes containing ligand artifacts, ions, or undefined ligands, and processing using Schrödinger’s LigPrep and PrepWizard, we utilized 250,633 protein-ligand systems.
SAIR
The Structurally Augmented IC50 Repository (SAIR) [12] is the largest publicly available dataset of protein-ligand 3D structures with binding affinity annotations, addressing the scarcity of high-quality experimental structures for deep learning applications. The original dataset contains 5,244,285 computationally generated structures across 1,048,857 unique protein-ligand systems from ChEMBL and BindingDB, with structures folded using the Boltz-1x model [13]. We applied stringent filtering criteria, retaining only PoseBusters-valid [14] complexes with negative AutoDock-Vina scores [15], confidence scores 0.8, interaction PTM 0.6, and IPTM 0.8, yielding 1,781,634 complexes. Further processing using Schrödinger’s LigPrep and PrepWizard resulted in 1,564,677 curated complexes with IC50 annotations.
BindingMOAD
BindingMOAD [16] is a comprehensive database developed over two decades (2001-2025), containing 41,409 protein-ligand complexes with affinity coverage for 15,223 complexes (37%) and 20,387 unique ligands. After preprocessing with Schrödinger’s LigPrep and PrepWizard, we obtained 33,286 complexes with diverse binding affinity annotations.
SPINDR
SPINDR [17] is a higher-fidelity dataset of protein–ligand complexes curated for interaction-aware modeling. The dataset emphasizes accurate binding site geometries and ligand poses, supporting physically consistent generative modeling. Complexes are thoroughly cleaned to focus on drug-like, non-covalent interactions using Schrödinger’s PrepWizard, with protein-ligand interactions annotated using ProLIF [18]. We employed all 35,627 provided complexes with partial binding affinity annotations.
HiQBind
HiQBind [19] is a curated dataset addressing structural artifacts in widely-used datasets, like PDBbind. Containing over 18,000 unique PDB entries and 30,000 protein-ligand complex structures, it matches binding free energies from BioLiP, Binding MOAD, and BindingDB with co-crystallized PDB complexes. The dataset employs strictly open-source curation tools with multiple quality control modules for steric clash detection, ligand structure fixing, protein completion, and hydrogen addition protocols. We utilized all 31,571 complexes with binding affinity annotations.
Schrodinger FEP+ dataset
The Schrodinger FEP+ dataset [20] comprises a large and diverse collection of protein-ligand complexes, each featuring congeneric series of small molecules with experimentally measured binding affinities (Kd, Ki, or IC50). Designed as a benchmark for assessing the accuracy and reproducibility of free energy perturbation (FEP) methods, the dataset emphasizes high-quality structural data, including X-ray structures and carefully curated protein and ligand preparations. It covers a broad chemical space, including challenging cases such as macrocycles, charge-changing transformations, and buried water displacement. This dataset is intended to support the development, validation, and comparison of computational methods for predicting relative binding affinities, providing a robust foundation for downstream applications in drug discovery and molecular design.
Data heterogeneity and quality control.
The use of heterogeneous, multi-fidelity data—including computationally generated protein–ligand complexes—is a deliberate and principled design choice supported by both foundational machine learning theory and recent domain-specific evidence. Transfer learning research has consistently shown that pretraining on diverse, even noisy, data improves downstream task performance: Yosinski et al. [21] demonstrated that transferred features outperform random initialization even across distant tasks, while Mahajan et al. [22] showed that pretraining on 3.5 billion weakly labeled images transfers remarkably well to clean benchmarks. In physics, McCabe et al. [23] showed that a single model pretrained on heterogeneous physics simulations outperforms task-specific baselines when finetuned on unseen systems.
Within drug discovery, the value of computationally generated structural data has been independently validated. Valsson et al. [24] demonstrated that augmenting training data with BindingNet complexes improved FEP benchmark performance (PCC , Kendall ). Most directly relevant, Hsu et al. [25] systematically showed that augmentation benefits depend critically on structural quality: high-confidence BindingNet structures substantially improve binding affinity prediction ( correlation between PCC and training set size), while moderate- and low-confidence structures provide negligible or negative returns. Co-folded structures (Boltz-1x) were shown to substitute effectively for experimental structures when filtered by confidence.
Our pipeline implements quality control measures that align with these findings: (1) SAIR complexes are filtered from 5M to 1.5M using stringent multi-criteria thresholds (PoseBusters-validity, Vina scores, Boltz-1x confidence); (2) BindingNet complexes are stratified by confidence, with only the high-confidence subset treated as higher-fidelity data; and (3) the progressive training paradigm ensures that Stage 2 finetuning on curated experimental data (SPINDR, HiQBind) overrides residual noise from lower-fidelity Stage 1 sources.
Dataset Statistics
@@
In Supplementary Section Datasets in Supplementary Fig. 10 we visualize the data generation framework that we established for this work. As outlined in Supplementary Section Datasets we used as protein-ligand complex data Plinder, BindingMOAD, SPINDR, HiQBind, SAIR, BindingNet, CrossDocked2020, KIBA-3D, Davis-3D, and Kinodata-3D. After downloading and aggregating the different data sources, we split proteins and ligands, if not already done, into separate files. We save ligands as SDF and proteins as PDB or CIF files. Afterwards, we filter the data and remove systems that classify the ligands either as ions, artifacts or not definable. Then, we use Schrodinger’s LigPrep to prepare the ligands and Schrodinger’s PrepWizard to prepare the protein-ligand complex (see Supplementary Section Datasets for more details). Successfully prepared systems are converted into our internal pocket-ligand representation encapsulated as PocketComplex instances for easy handling and data management. We retrieve pockets by cutting around the reference ligands with a cutoff of 7A and, if necessary, trim the pocket to the max size of 800 atoms. Every PocketComplex instance is annotated with comprehensive metadata spanning affinity values (if available), GNINA scores, interaction fingerprints, and selected RDKit features, like molecular weight, logP, TPSA, number of hydrogen acceptors and donors, number of rotatable bonds, number of aliphatic rings and many more. Finally, for all datasets we calculate distribution statistics comprising atom, charge, and bond type distributions as well as bond angles, bond lengths and dihedral angles distributions. Importantly, all datasets are splitted following the Plinder train, validation and test splits.
@@
In Supplementary Fig. 11, we compared the chemical composition of all protein–ligand datasets used in this work by analyzing the distributions of ligand atom types, protein atom types, and protein residue types. The aggregated distributions (top row) reveal the overall prevalence of each type, while the per-dataset heatmaps (bottom row) illustrate how these distributions vary between datasets. For atom types, a logarithmic scale was used to emphasize less frequent elements. Only the 20 standard amino acids were considered for residue type comparisons, and hydrogens were excluded from atom type analyses to improve interpretability.
@@
To provide a comprehensive overview of the affinity data, Supplementary Fig. 12 summarizes both the distribution and coverage of affinity types across all datasets. The ECDFs display the range and distribution of affinity values for each affinity type, highlighting differences in spread and modality. The overall affinity type counts, presented on a logarithmic scale, reveal a pronounced imbalance, with IC50 being the most abundant. Dataset-specific distributions are visualized using horizontal violin plots, while horizontal bar plots summarize the number of measurements for each affinity type within each dataset. The majority of affinity values originate from the SAIR and BindingNet datasets. Notably, BindingNet also provides Ki measurements, while HiQBind and BindingMOAD offer a balanced representation of IC50, Ki, and Kd values. HiQBind is the only dataset containing EC50 measurements. In contrast, Davis-3D is exclusively annotated with Kd values. This overview facilitates the identification of data biases and gaps that may influence downstream analyses.
@@
To systematically assess the chemical diversity of ligands across datasets, we analyzed the distributions of several continuous molecular descriptors in Supplementary Fig. 13, including molecular weight, logP, molar refractivity, and topological polar surface area (TPSA). These features capture key aspects of ligand size, hydrophobicity, polarizability, and polarity, which are relevant for molecular recognition and drug-likeness. The results reveal pronounced differences between datasets: for example, BindingNet and SAIR contain ligands with a broader range of molecular weights and higher TPSA values, while datasets such as HiQBind and BindingMOAD exhibit more constrained distributions. These trends reflect the varying selection criteria and source domains of the datasets, and highlight the importance of considering chemical diversity in benchmarking and model development.
@@
We further characterized the datasets by comparing the distributions of discrete ligand features in Supplementary Fig. 14, including the number of hydrogen bond acceptors and donors, rotatable bonds, ring systems, and chiral centers. These properties provide insight into molecular complexity, flexibility, and the presence of functional groups relevant for binding interactions. The analysis demonstrates that datasets such as BindingNet and SAIR encompass ligands with higher numbers of rotatable bonds and greater ring system diversity, whereas datasets like Davis-3D and KIBA-3D are more restricted in these aspects. Notably, the number of chiral centers and structural alerts also varies substantially, underscoring differences in stereochemical complexity and potential reactivity. Together, these comparisons elucidate the distinct chemical spaces sampled by each dataset and inform the interpretation of downstream modeling results.
KIBA-3D
@@
The KIBA dataset was originally introduced by Tang et al. [26] as a large-scale benchmark for kinase–ligand binding affinity prediction, integrating heterogeneous bioactivity readouts into a standardized KIBA score. The original dataset contained chemical compounds, kinase targets and over interaction records. While comprehensive, the original KIBA dataset was highly scattered, due to an imbalance between the number of ligand-kinase pairs and associated measurements. To improve the density and ensure more reliable evaluation, subsequent studies discarded the ligands and targets with less than 10 interaction records. This procedure reduced the dataset to ligands and kinase targets. For this work, we prepared the 3D extension of KIBA (KIBA-3D) based on kinase targets with experimentally solved structures deposited in the PDB. Supplementary Fig. 15 shows examples of our docking results, illustrating that Glide could reproduce near-native binding poses for selected kinase inhibitors with conserved bidentate hydrogen bond pattern in the hinge region (for example, PDBID:2XB7-CHEMBL461876, glide gscore , PDBID:6LVM-CHEMBL1968590, glide gscore ). The KIBA-3D dataset serves as a large-scale, structure-consistent benchmark for evaluating generative and predictive models.
Target protein structure selection for KIBA-3D